# Multi-Task Self-Supervised Learning for Generalizable IQ Representations: Toward Foundation Models for AI-Native 6G

### v1

@@ I INTRODUCTION
To address this challenge, we introduce an SSL framework designed specifically for unprocessed multi-channel MIMO IQ data.
This framework supports efficient multi-task learning across diverse wireless tasks by preserving the unique signal characteristics necessary for each task. The key contributions of this work are summarized as follows:
@@
  • We present, to the best of our knowledge, the first foundational model for wireless communication signals that operates directly on raw IQ data. The model supports a variety of tasks, including modulation classification, angle-of-arrival estimation, RF fingerprinting, and beam prediction, demonstrating multi-task learning capabilities across diverse wireless tasks. Our architecture employs a lightweight ShuffleNetV2 (0.5×) backbone, enabling efficient feature extraction and real-time inference while maintaining a compact model footprint suitable for resource-constrained deployment.
@@ II-B Multi-Task Learning in Wireless Communication
Multi-task learning (MTL) has been investigated in wireless communication to enable the joint modeling of related tasks. In [26], a fully supervised MTL model was proposed that operates on raw IQ data for simultaneous modulation and protocol classification, utilizing a shared encoder with task-specific output heads. However, their approach remains fully supervised and does not incorporate self-supervised learning principles. While self-supervised multi-task learning has been explored using preprocessed representations such as channel state information (CSI) or channel impulse responses (CIR) [16], efforts to extend SSL directly to raw IQ signals are still limited.
@@ II-C Foundational Models for Wireless Communication
Inspired by the success of large-scale pretraining in vision and language, foundation models are gaining traction in wireless communication, particularly in the context of emerging 6G and integrated sensing and communication (ISAC) systems. CSI-CLIP[18] introduces a contrastive learning framework that treats CSI and CIR as aligned modalities to pretrain a MIMO foundation model, supporting downstream tasks such as positioning and beam management. WirelessGPT[16] leverages unsupervised pretraining on large-scale datasets to learn universal representations, enabling applications ranging from channel prediction to human activity recognition. WavesFM[17] adopts a ViT-based architecture with task-specific heads and LoRA, using masked modeling on spectrograms and CSI to support a wide range of communication and sensing tasks. While these models demonstrate versatility, they typically rely on structured inputs such as CSI, CIR, or spectrograms, which require extensive preprocessing. In contrast, foundational models that operate directly on raw IQ data remain largely unexplored. Our work addresses this gap by proposing a raw-IQ-based FM that enables multi-task learning without handcrafted features and generalizes to unseen downstream tasks, offering improved adaptability across domains.
@@ V-A Encoder Architecture and Experimental Setup
We employ the ShuffleNetV2 (0.5×) architecture as the encoder backbone due to its lightweight structure and suitability for deployment on edge devices. Our implementation, based on the official PyTorch release, contains approximately 1.3 million trainable parameters and leverages depthwise separable and grouped convolutions to optimize for inference efficiency. The encoder is followed by a task-specific projection head during contrastive pretraining, and lightweight classification heads are attached during the fine-tuning phase.
@@ V-D Evaluation Results on Multiple Datasets
To ensure input compatibility, RF fingerprinting and DeepBeam signals are segmented into 256-sample frames with missing channels padded by zeros, while the single-channel 128-sample RML2016 samples are processed via the encoder’s global pooling layer. For all experiments, the encoder remains frozen, and two lightweight adaptation strategies are evaluated: (i) linear probing, which appends a linear classifier on top of the frozen encoder, and (ii) Low-Rank Adaptation, which introduces approximately 84K trainable parameters into the convolutional layers. In contrast, the fully supervised baseline requires training all 1.3M parameters of the ShuffleNetV2 (0.5×) backbone.

### v2

@@ I INTRODUCTION
To address this, we introduce an SSL framework designed specifically for unprocessed multi-channel MIMO IQ data. This framework supports efficient multi-task learning across diverse wireless tasks by preserving the unique signal characteristics necessary for each task. The key contributions of this work are summarized as follows:
@@
  • We present, to the best of our knowledge, the first foundational model for wireless communication signals that operates directly on raw IQ data. The model supports a variety of tasks, including modulation and angle-of-arrival classification, RF fingerprinting, and beam prediction, demonstrating multi-task learning capabilities across diverse wireless tasks. Our architecture employs a lightweight ShuffleNetV2 (0.5×) backbone, enabling efficient feature extraction and real-time inference while maintaining a compact model footprint suitable for resource-constrained deployment.
@@ II-B Multi-Task Learning in Wireless Communication
Multi-task learning (MTL) has been investigated in wireless communication to enable the joint modeling of related tasks. In [26], a fully supervised MTL model was proposed that operates on raw IQ data for simultaneous modulation and protocol classification, utilizing a shared encoder with task-specific output heads. However, their approach remains fully supervised and does not incorporate self-supervised learning principles. While self-supervised multi-task learning has been explored using preprocessed representations such as CSI or CIR [16], efforts to extend SSL directly to raw IQ signals are still limited.
@@ II-C Foundational Models for Wireless Communication
Inspired by the success of large-scale pretraining in vision and language, foundation models are gaining traction in wireless communication, particularly in the context of emerging 6G and integrated sensing and communication (ISAC) systems. WavesFM [17] builds upon earlier FM work presented by the same authors in [28, 29], where they introduced and subsequently improved the first wireless FM leveraging SSL to extract generalizable representations from RF spectrograms. Their work demonstrated effective performance in multi-task scenarios, including human activity sensing and spectrogram segmentation, using a single shared FM. WavesFM extends these approaches further by adopting a ViT-based architecture with task-specific heads and LoRA, employing masked modeling techniques on spectrograms and CSI, trained on real-world data to support a broader spectrum of communication, sensing, and localization tasks.
Wireless channel representation learning has also gained attention in recent literature. CSI-CLIP [18] introduced a contrastive learning method, treating CSI and CIR as aligned modalities to pretrain a wireless channel foundation model, facilitating tasks like positioning, beam management, and line-of-sight (LOS) versus non-line-of-sight (NLOS) identification. Similarly, WirelessGPT [16] employed a masked reconstruction strategy to learn robust representations from extensive simulated datasets, enabling effective performance in channel prediction and human activity recognition tasks. Both methods primarily rely on simulated data during pre-training and fine-tuning phases.
While these models demonstrate versatility, they typically rely on structured inputs such as CSI, CIR, or spectrograms, which require extensive preprocessing. In contrast, foundational models that operate directly on raw IQ data remain largely unexplored. Our work addresses this gap by proposing a raw-IQ-based FM that enables multi-task learning without handcrafted features and generalizes to unseen downstream tasks, offering improved adaptability across domains
@@ V-A Encoder Architecture and Experimental Setup
We employ the ShuffleNetV2 (0.5×) architecture as the encoder backbone due to its lightweight structure and suitability for deployment on edge devices. Our implementation, based on the official PyTorch release, contains approximately 341k trainable parameters and leverages depthwise separable and grouped convolutions to optimize for inference efficiency. The encoder is followed by a task-specific projection head during contrastive pretraining, and lightweight classification heads are attached during the fine-tuning phase.
@@ V-D Evaluation Results on Multiple Datasets
To ensure input compatibility, RF fingerprinting and DeepBeam signals are segmented into 256-sample frames with missing channels padded by zeros, while the single-channel 128-sample RML2016 samples are processed via the encoder’s global pooling layer. For all experiments, the encoder remains frozen, and two lightweight adaptation strategies are evaluated: (i) linear probing, which appends a linear classifier on top of the frozen encoder, and (ii) Low-Rank Adaptation, which introduces approximately 84K trainable parameters into the convolutional layers. In contrast, the fully supervised baseline trains the entire ShuffleNetV2 (0.5×) backbone comprising 341.8K parameters, excluding the final task-specific linear layer whose size varies with the number of output classes.